# BotArtist: Generic approach for bot detection in Twitter via semi-automatic machine learning pipeline

### v3

@@ Introduction
The results, as presented in table: 4, indicate that the most generalizable models are BotArtist, Botometer, and SGBot. BotArtist demonstrated superior performance over existing methods in terms of both F1-total and average scores, surpassing almost 10% existing methods in total. These results confirm that fine-tuned models which are based on the limited set of features are capable of capturing the general differences between normal and bot accounts.
Furthermore, it’s noteworthy that the BotArtist model outperformed existing methods in both comparison scenarios. Due to the high volume of experiments, we do not provide any information according to the selected features and final model configuration for each evaluation scenario. These data will be shared and described in detail, in a GitHub repository after the paper acceptance.
@@
Explainability
Besides the creation of the high-performance model used for bot detection, we aim to shed light on the model decision itself. For this purpose we employ SHapley Additive exPlanations (SHAP) values (Lundberg and Lee 2017) which offer various benefits. One of the most significant advantages of SHAP values is that they are not tied to any particular type of machine learning model and are therefore model-agnostic. Furthermore, SHAP values display local accuracy, and consistency, which are characteristics that are not found concurrently in other methods.
In this section we explain the BotArtist model trained over multiple datasets. Figure: 2 presents the 20 most important features and their impact on the final decision. The features are sorted in descending order of importance. Based on this order, we observe that the age of the account and the level of social activity are the most critical characteristics of our model, since the creation date of the account.
The presented figure showcases the values of the features, ranging from low (indicated by the blue color), to high (indicated by the red color). Simultaneously, the x-axis provides information on the impact of each feature value on the model decision. Based on the presented results, it is crucial to highlight that the SHAP explanation technique provides insights that align with past related works. For instance, we find that bot accounts tend to have a significantly lower age of the account and lower number of followers combined with a higher number of followers by age of the account. Based on this information we can assume that bot accounts are not so popular in terms of total followers in comparison with normal users, but the growth of the profile followers is higher in comparison with normal users. This assumption agrees with the original target of the bot account creators to achieve a higher audience during the short period.
Conclusions and Future Work
In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach. We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features. Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets. As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in total F1-score in multiple dataset comparisons and a 6.5% increase in separated dataset comparisons. Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach. Upon acceptance of this paper, we will make the developed SAMLP pipeline available for further research, along with the BotArtist model, which can label existing profile data. One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features. Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social media.

### v4

@@ 5.1 Models comparison
The results, as presented in Table 4, indicate that the most generalizable models are BotArtist, Botometer, and SGBot. BotArtist demonstrates superior performance over existing methods in terms of both total and average scores, surpassing almost 10% the existing methods in total. These results confirm that fine-tuned models which are based on the limited set of features are capable of capturing the general differences between normal and bot accounts.
6 Conclusions and Future Work
In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach. We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features. Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets. As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in the total F1 score in comparisons of multiple data sets and a 6. 5% increase in comparisons of separated data sets. Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach (available at the repository). One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features. Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social networks.
For further research, we provide one of the largest labeled datasets containing user profiles correlated with another publicly available dataset of the 127M user posts. Such large labeled data will allow the development of state-of-the-art LLM models to detect bot-generated content in the post-Twitter API era.
Appendix A Appendix
A.1 Model benchmarks
As mentioned in section 5 we compare BotArtist implementation with the existing 35 bot detection methods. We replicate the experiments of the model benchmarks described in (Feng et al., 2022a) where the authors provide the implementation of the 34 models. The Botometer (Yang et al., 2022) is an available model on-line that does not require additional training. Furthermore, the authors provide the original dataset split for training, validation, and testing portions, which allows the replication of the experiments with unbiased configurations. We manage to re-train each model (except the Botometer) over the training/validation set and evaluate their performance over the test set, as provided by (Feng et al., 2022a) (see Table 3).
We differentiate from the original paper in the generic case scenario, as we combine the training/validation portions of the nine datasets into a single train portion where the selected top-performance models were trained over the combination of the datasets to evaluate the generalization of the methods. To reproduce the experiments, the repository  [^4] GitHub repository: TBD contains all the necessary scripts and information. The designed experiments were executed on a Dell server machine with two AMD EPYC 7543 CPUs (32 cores/64 threads), 512GB DDR4 RAM (16X32 GB) and nVidia A30 GPUs (24 GB VRAM), with a total execution time around 30 days.
A.2 Created dataset
Table 5: Set of hashtags used in our data collection query written in Russian, Ukrainian, and English provided with translation for non-English hashtags.
#Ukraine, #Ukraina, #ukraina, #Украина(Ukraine), #Украине(Ukraine), #PrayForUkraine, #UkraineRussie, #StandWithUkraine, #StandWithUkraineNOW, #RussiaUkraineConflict, #RussiaUkraineCrisis, #RussiaInvadedUkraine, #WWIII, #worldwar3, #Война(War), #BlockPutinWallets, #UkraineRussiaWar, #Putin, #Russia, #Россия(Russia), #StopPutin, #StopRussianAggression, #StopRussia, #Ukraine_Russia, #Russian_Ukrainian, #SWIFT, #NATO, #FuckPutin, #solidarityWithUkraine, #PutinWarCriminal, #PutinHitler, #BoycottRussia, #with_russia, #FUCK_NATO, #ЯпротивВойны(I’m against war), #StopNazism #myfriendPutin #UnitedAgainstUkraine #StopWar #ВпередРоссия(Go Russia), #ЯМыРоссия(I/we Russia), #ВеликаяРоссия(Great Russia), #Путинмойпрезидент(Putin is my president), #россиявперед(Go Russia), #россиявперёд(Go Russia), #ПутинНашПрезидент(Putin is our president), #ЗаПутина(For Putin), #ПутинВведиВойска(Putin send the troops), #СЛАВАРОССИИ(Glory to Russia), #СЛАВАВДВ(Glory to Russian Airborne Forces)
The provided 10M Twitter profiles dataset was retrieved through the Twitter API, while still available to the research community. The collection of the user profiles was initiated at the beginning of the 2022 Russo-Ukrainian War on 24 February 2022. We identify most of the popular hashtags based on the Ukrainian, Russian, and English hashtags during the initial days of invasion (see Table 5). Based on the collected hashtags we initiate the collection of the respective user profiles using them in their posts. Additionally, by using Twitter API we collect data from previous periods (23 February 2022) to capture some profiles and their activity 24 hours before the invasion. Furthermore, we collect on a daily basis the user profiles for the 16-month period until June 23, 2023.
To share the collected user profiles and not violate user privacy and rights of the removal of their information from online social media, we provide the dataset in the processed form of CSV files where only numerical features are presented without any uuid information such as username, screen_name, description, URLs, user_id and profile_image_url. To hide the private user data while providing as much useful information for further usage in the case of ML/AI applications, we compute as many features as possible to maximize the information provided by the dataset (see Table 2).
Additionally, based on the described BotArtist model we predict the collected user profiles and reveal which profile belongs to a human or bot account, based on BotArtist in combination with the model probability. Furthermore, we contact the original authors of the existing anonymized textual dataset over the same topic of Russo-Ukrainian War (Shevtsov et al., 2022a). We provide identical non-reversible anonymized uids to keep the connection between the user profile information and their tweets or user mentions over the 127.275.386 tweets. We believe that the combination of the provided labels and the large volume of textual data in the new era of Large Language Models will help the research community to develop state-of-the-art solutions for bot content detection. For this purpose the provided Zenodo dataset  [^5] Zenodo data repository: TBD and the metadata  [^6] GitHub repository: TBD are shared over the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 DEED)  [^7] Dataset license: https://creativecommons.org/licenses/by-nc-sa/4.0/. Shared data will not be maintained due to updated Twitter API subscriptions and will be provided only in the current public version of Zenodo repository in form of CSV file (see data sample provided in Table 6 ).
Table 6: Sample of the provided data with the entire feature set.
age description_urls description_mentions description_hashtags screen_name_sim statuses statuses_by_age following following_by_age followers followers_by_age listed listed_by_age 786.0 0 0 0 0.2 1215 1.545801526717557 258 0.3282442748091603 80 0.1017811704834605 0 0.0 3410.0 0 0 0 0.4117647058823529 178614 52.37947214076247 95 0.0278592375366568 1642 0.481524926686217 41 0.0120234604105571 4266.0 2 0 0 0.8571428571428571 207601 48.66408813877168 8 0.0018752930145335 3911749 916.9594467885609 1895 0.4442100328176278 2083.0 0 0 0 0.6923076923076923 1271 0.6101776284205472 138 0.0662506000960153 30 0.0144023043686989 4 0.0019203072491598 5224.0 0 3 0 0.4166666666666667 27539 5.271630934150076 1394 0.2668453292496171 2006632 384.1179173047473 8880 1.6998468606431854 870.0 0 0 0 0.3076923076923077 2331 2.679310344827586 325 0.3735632183908046 152 0.1747126436781609 0 0.0 1239.0 1 2 0 1.0 33541 27.07102502017756 5 0.0040355125100887 2514015 2029.0677966101696 $2$ 1.7933817594834545 1328.0 0 0 0 0.0 10589 7.973644578313253 318 0.2394578313253012 193 0.1453313253012048 6 0.0045180722891566 1127.0 1 0 1 0.2666666666666666 1728 1.5332741792369122 44 0.0390417036379769 15908 14.115350488021296 213 0.188997338065661 2202.0 1 0 0 0.4 5059 2.297456857402361 110 0.0499545867393278 8959 4.068574023614896 6 0.0027247956403269 13.0 0 0 0 0.2105263157894736 746 57.38461538461539 0 0.0 49506 3808.153846153846 267 20.53846153846154 1288.0 0 0 0 1.0 237937 184.7336956521739 4758 3.694099378881988 7546 5.858695652173913 46 0.0357142857142857 4323.0 1 0 0 0.6153846153846154 39331 9.098080037011334 610 0.1411057136247976 26450 6.118436271108027 474 0.1096460791117279 3422.0 0 0 0 0.7142857142857143 182898 53.44769140853303 5271 1.540327293980129 5131 1.4994155464640562 5 0.0014611338398597 899.0 0 0 3 0.5714285714285714 8028 8.92992213570634 4864 5.410456062291435 3862 4.295884315906563 2 0.0022246941045606 925.0 0 0 0 0.8 401921 434.5091891891892 6203 6.705945945945946 9057 9.791351351351352 1 0.001081081081081 781.0 0 0 0 0.1333333333333333 17805 22.797695262483995 13313 17.0460947503201 17091 21.88348271446863 10 0.0128040973111395 4628.0 0 0 0 0.8181818181818182 7642 1.6512532411408816 1670 0.3608470181503889 87651 18.939282627484875 358 0.0773552290406223 4771.0 0 0 0 0.8181818181818182 27719 5.809893104171033 2558 0.5361559421504926 3283 0.6881156990148816 120 0.0251519597568643 profile_image description_upper_len description_lower_len description_digit_len description_spec_len description_upper_pcnt description_lower_pcnt description_digit_pcnt description_spec_pcnt description_len screen_name_upper_len screen_name_lower_len screen_name_digit_len 1 1 91 0 24 0.0086206896551724 0.7844827586206896 0.0 0.2068965517241379 116 0 12 0 1 0 0 0 153 0.0 0.0 0.0 1.0 153 0 13 1 1 10 34 0 68 0.0892857142857142 0.3035714285714285 0.0 0.6071428571428571 112 0 9 0 1 5 94 0 29 0.0390625 0.734375 0.0 0.2265625 128 3 12 0 1 22 93 0 38 0.1437908496732026 0.6078431372549019 0.0 0.2483660130718954 153 0 10 0 1 4 33 0 52 0.0449438202247191 0.3707865168539326 0.0 0.5842696629213483 89 2 2 3 1 12 72 1 36 0.0991735537190082 0.5950413223140496 0.0082644628099173 0.2975206611570248 121 1 9 0 1 0 35 0 7 0.0 0.8333333333333334 0.0 0.1666666666666666 42 0 10 0 1 26 75 2 30 0.1954887218045112 0.5639097744360902 0.0150375939849624 0.225563909774436 133 2 5 0 1 3 111 0 28 0.0211267605633802 0.7816901408450704 0.0 0.1971830985915492 142 0 11 0 1 43 18 6 13 0.5375 0.225 0.075 0.1625 80 2 0 5 1 3 85 0 23 0.027027027027027 0.7657657657657657 0.0 0.2072072072072072 111 1 7 3 1 14 102 4 23 0.0979020979020979 0.7132867132867133 0.0279720279720279 0.1608391608391608 143 0 14 0 1 3 47 0 17 0.044776119402985 0.7014925373134329 0.0 0.2537313432835821 67 1 5 2 1 4 69 0 28 0.0396039603960396 0.6831683168316832 0.0 0.2772277227722772 101 0 10 1 1 5 35 0 10 0.1 0.7 0.0 0.2 50 0 9 1 1 11 55 0 31 0.1134020618556701 0.5670103092783505 0.0 0.3195876288659793 97 2 9 0 1 14 88 2 22 0.1111111111111111 0.6984126984126984 0.0158730158730158 0.1746031746031746 126 0 14 0 1 15 105 0 39 0.0943396226415094 0.660377358490566 0.0 0.2452830188679245 159 0 10 0 screen_name_spec_len screen_name_upper_pcnt screen_name_lower_pcnt screen_name_digit_pcnt screen_name_spec_pcnt screen_name_len name_upper_len name_lower_len name_digit_len name_spec_len name_upper_pcnt name_lower_pcnt name_digit_pcnt 0 0.0 1.0 0.0 0.0 12 2 8 0 1 0.1818181818181818 0.7272727272727273 0.0 0 0.0 0.9285714285714286 0.0714285714285714 0.0 14 4 10 0 0 0.2857142857142857 0.7142857142857143 0.0 0 0.0 1.0 0.0 0.0 9 5 4 0 1 0.5 0.4 0.0 0 0.2 0.8 0.0 0.0 15 2 10 0 1 0.1538461538461538 0.7692307692307693 0.0 0 0.0 1.0 0.0 0.0 10 3 15 0 2 0.15 0.75 0.0 2 0.2222222222222222 0.2222222222222222 0.3333333333333333 0.2222222222222222 9 1 3 0 5 0.1111111111111111 0.3333333333333333 0.0 0 0.1 0.9 0.0 0.0 10 1 9 0 0 0.1 0.9 0.0 1 0.0 0.9090909090909092 0.0 0.0909090909090909 11 8 5 0 2 0.5333333333333333 0.3333333333333333 0.0 0 0.2857142857142857 0.7142857142857143 0.0 0.0 7 3 12 0 6 0.1428571428571428 0.5714285714285714 0.0 1 0.0 0.9166666666666666 0.0 0.0833333333333333 12 1 3 0 0 0.25 0.75 0.0 0 0.2857142857142857 0.0 0.7142857142857143 0.0 7 15 8 5 9 0.4054054054054054 0.2162162162162162 0.1351351351351351 0 0.0909090909090909 0.6363636363636364 0.2727272727272727 0.0 11 1 7 2 0 0.1 0.7 0.2 1 0.0 0.9333333333333332 0.0 0.0666666666666666 15 3 19 0 3 0.12 0.76 0.0 0 0.125 0.625 0.25 0.0 8 1 5 0 0 0.1666666666666666 0.8333333333333334 0.0 1 0.0 0.8333333333333334 0.0833333333333333 0.0833333333333333 12 0 14 0 1 0.0 0.9333333333333332 0.0 0 0.0 0.9 0.1 0.0 10 0 9 0 1 0.0 0.9 0.0 4 0.1333333333333333 0.6 0.0 0.2666666666666666 15 1 6 0 6 0.0769230769230769 0.4615384615384615 0.0 0 0.0 1.0 0.0 0.0 14 2 14 0 1 0.1176470588235294 0.8235294117647058 0.0 0 0.0 1.0 0.0 0.0 10 2 11 0 1 0.1428571428571428 0.7857142857142857 0.0 name_spec_pcnt name_len screen_name_entropy name_entropy has_location profile_url total_urls foll_friends verified protected uid label proba 0.0909090909090909 11 2.138333059508028 2.2718685126965625 1 1 1 0.3100775193798449 0 0 uidb46b… human 68 0.0 14 2.4410152780267027 2.4410152780267023 1 0 0 17.28421052631579 0 0 uida3ea… human 99 0.1 10 1.7351264569629226 2.302585092994045 1 1 3 488968.625 1 0 uidde3a… human 99 0.0769230769230769 13 2.5232109529528914 2.351673301904631 0 0 0 0.217391304347826 0 0 uid9dcc… bot 54 0.1 20 1.4708084763221112 2.4843668399738017 1 1 1 1439.477761836442 1 0 uid2991… human 99 0.5555555555555556 9 2.0431918705451206 2.19722457733622 1 0 0 0.4676923076923077 0 0 uida01b… human 95 0.0 10 2.0253262207700677 2.0253262207700677 1 1 2 502803.0 1 0 uid12f2… human 99 0.1333333333333333 15 1.9722469794234416 2.5232109529528914 1 1 1 0.6069182389937107 0 0 uidce17… human 93 0.2857142857142857 21 1.5498260458782016 2.532591160801121 1 1 2 361.54545454545456 0 0 uid8d37… human 91 0.0 4 2.253857589601352 1.3862943611198906 1 1 2 81.44545454545455 0 0 uid569e… human 93 0.2432432432432432 37 1.277034259466139 2.9147497445394843 1 0 0 0.0 0 0 uid5955… bot 92 0.0 10 2.0198149924929463 2.0253262207700677 0 0 0 1.585960487599832 0 0 uid5ef3… human 99 0.12 25 2.026229622995292 2.310717508000889 1 1 2 43.36065573770492 0 0 uidac7a… human 99 0.0 6 1.9061547465398496 1.5607104090414063 0 0 0 0.9734395750332006 0 0 uid059f… human 99 0.0666666666666666 15 2.2538575896013526 2.430791328878232 1 0 0 0.7939967105263158 0 0 uid55ca… human 94 0.1 10 2.1639556568820564 2.1639556568820564 0 0 0 1.460099951636305 0 0 uid9437… human 99 0.4615384615384615 13 2.2459520807289133 1.8446214763655029 1 1 1 1.2837827687223016 0 0 uid10f1… human 95 0.0588235294117647 17 2.106577333232116 2.3131536787336344 1 1 1 52.48562874251497 1 0 uid975d… human 99 0.0714285714285714 14 2.1639556568820564 2.4410152780267027 0 0 0 1.2834245504300237 0 0 uid4d4c… human 99
@@ A.2 Created dataset
A.3 Explainability
This study aims to create a high performance model that can be used for bot detection while shedding light on the decision of the model itself. For this purpose, we employ SHapley Additive exPlanations (SHAP) values (Lundberg and Lee, 2017) which offer various benefits. One of the most significant advantages of SHAP values is that they are not tied to any particular type of machine learning model and, therefore, are model-agnostic. Furthermore, SHAP values display local accuracy and consistency and cannot be found concurrently in other methods.
In this section, we explain the BotArtist model trained over multiple datasets. Figure: 2 presents the 20 most important features and their impact on the final decision. The features are sorted in descending order of importance. Based on this order, we observe that the age of the account and the level of social activity are the most critical characteristics of our model, since the creation date of the account.
The presented figure showcases the values of the features, ranging from low (indicated by the blue color) to high (indicated by the red color). The x-axis provides information on the impact of each feature value on the model decision. Based on the results presented, it is crucial to note that the SHAP explanation technique provides insights that align with past related work. For instance, we reveal that bot accounts tend to have a significantly lower age of the account and a lower number of followers combined with a higher number of followers by age of the account. Based on this information, we can assume that bot accounts are not so popular in terms of total followers in comparison with normal users, while the growth of the profile followers is higher in comparison with normal users. This assumption agrees with the original target of the bot account creators to achieve a higher audience during the short period.